# Convergence of Decentralized Stochastic Subgradient-based Methods for Nonsmooth Nonconvex functions

### v3

@@ 2.2 Mixing matrix
Mixing matrix ${\bm{W}}=[{\bm{W}}(i,j)]\in\mathbb{R}^{d\times d}$ determines the structure of the communication network and plays an important role in averaging the local information of neighbor agents. Generally, we assume the mixing matrix ${\bm{W}}$ satisfies the following properties, which are standard in the literature [52, 53, 9].
@@
With a given graph $\mathtt{G}$, there are various approaches to choose its corresponding mixing matrix, such as Laplacian-based constant edge weight matrix [52] and Metropolis weight matrix [54]. We refer interested readers to [55, 56] for more details about how to choose the mixing matrix.
Based on [57, Perron-Frobenius Theorem], we derive a direct corollary, which illustrates the fundamental spectral property of mixing matrix ${\bm{W}}$.
@@
For any mixing matrix ${\bm{W}}\in\mathbb{R}^{d\times d}$ that corresponds to a connected graph $\mathtt{G}$, all the eigenvalues of ${\bm{W}}$ lie in $(-1,1]$, and ${\bm{W}}$ has a single eigenvalue at $1$ that admits ${\bm{1}_{d}}$ as its eigenvector.
@@ 2.3 Nonsmooth analysis and conservative field
[33] For any given locally Lipschitz continuous function $f:\mathbb{R}^{n}\to\mathbb{R}$, and any ${\bm{x}}\in\mathbb{R}^{n}$, the generalized directional derivative of $f$ at ${\bm{x}}$ along the direction ${\bm{d}}\in\mathbb{R}^{n}$, is defined as
@@
[33] We say that $f$ is Clarke regular at ${\bm{x}}\in\mathbb{R}^{n}$, if for every direction ${\bm{d}}\in\mathbb{R}^{n}$, the one-sided directional derivative
@@
It is remarkable to note that any conservative field is locally bounded [36, Remark 3]. We now introduce the definition of the path-differentiable function corresponding to the conservative field.
@@
As illustrated in [36, Corollary 1], for any given path-differentiable function $f$, $\partial f$ is a conservative field for $f$. Therefore, the concept of the conservative field can be regarded as an extension of the Clarke subdifferential. More importantly, when conservative field $D_{f}({\bm{x}})$ is convex-valued, ${\bm{0}}\in\partial f({\bm{x}})$ implies that $0\in D_{f}({\bm{x}})$.
In fact, the class of path-differentiable functions is general enough to cover the most objective functions in real-world problems. The well-known Clarke regular functions mentioned in [33] and semi-algebraic functions mentioned in [58] are all path-differentiable functions. In [37, 36], the authors identify definable functions (whose graphs are definable in an o-minimal structure) as an important class of path-differentiable functions. Actually, most activation functions and loss functions in deep neural networks are definable functions, including sigmoid, softplus, ReLU, leaky ReLU, hinge loss, etc.
Furthermore, it is important to highlight that definability is preserved under finite summation and composition [37]. Consequently, for any neural network constructed by definable blocks, its loss function is definable, thus it is a path-differentiable function. Additionally, it is worth noting that the Clarke subdifferential of definable functions is itself definable [36]. Hence, in the case of a neural network constructed by definable blocks, the conservative field corresponding to the AD algorithms is also definable. The following proposition demonstrates that the definability of both $f$ and its conservative field $D_{f}$ leads to the nonsmooth Morse–Sard property [59].
Proposition 2.7 (Theorem 5 in [36]).
@@ 2.4 Differential inclusion and stochastic subgradient methods
Lemma 2.13 (Extension of Proposition 1.3 in [45]).
@@
Before exhibiting the results on uniform boundedness and convergence of (2.3), we provide an assumption about differential inclusion (2.2), which is commonly used in the analysis of ODE-based approaches [46, 47, 48].
@@
Lemma 2.15 (Theorem 3.6 in [43]).
@@
Lemma 2.16 (Summary of Theorem 3.6, Proposition 3.27 in [45]).
@@
Except for Lemma 2.16, similar results under slightly different conditions can be found in [47, 37, 48], while some recent works [60, 44] focus on analyzing the convergence of (2.3) under relaxed conditions. Interested readers could refer to those works for details.
@@ 3 A General Framework for Decentralized Stochastic Subgradient-based Methods
In this section, we aim to establish the convergence properties of proposed framework (DSM) under two noise settings. To this end, it is crucial to ensure the uniform boundedness of iterates $\{{\bm{Z}}_{k}\}$. A large amount of literature [38, 41, 37, 42] regards this condition as a prior assumption when analyzing the convergence of SGD-type algorithms in a single-agent setting. Nevertheless, more recent works [61, 49, 62, 63, 64, 43] focus on rigorously establishing the uniform boundedness of the iterative sequence.
@@ 3.1 Basic assumptions and main results
Assumption 3.1(1) allows for a flexible choice of the step-size, while Assumption 3.1(2) is reasonable and mild, which is commonly employed in various existing works [36, 37, 43, 40, 38].
Assumption 3.1(3) clarifies the Lyapunov properties of a differential inclusion with respect to $-\Phi$, which is standard in the literature [45, 65, 66, 37, 36, 63] and can be satisfied by most appropriately selected set-valued mappings. Since each $\Phi_{i}$ is graph-closed and locally bounded, the convex-valued set-valued mapping $\Phi$ is also graph-closed and locally bounded. Then, the coercivity of $\psi$ in Assumption 3.1(3) indicates the compactness of any level set $\mathcal{L}_{r}:=\{x\in\mathbb{R}^{n}:\psi(x)\leq r\}$.
Assumption 3.1(4) is similar to the Weak Sard’s condition in [37], which stipulates that the set of values of $\psi$ at points outside the stable set is dense in $\mathbb{R}$. As demonstrated in [36], Assumption 3.1(4) holds when $f$ is a definable function with the selection $\Phi:=\partial f$ and $\psi:=f$.
@@
[…]
@@
The following lemma from [53, Lemma 3.1] concerns the limit property of a convolution-like scalar sequence.
@@
Let $\{\gamma_{k}\}$ be a sequence of real numbers. If $10<a<$ 1 and $\lim_{k\rightarrow\infty}\gamma_{k}=\gamma$, then it holds that $\lim_{k\rightarrow\infty}\sum_{\ell=0}^{k}a^{k-\ell}\gamma_{\ell}=\frac{\gamma}{1-a}$.
@@ 3.2 Convergence with random reshuffling
Suppose Assumption 3.1 and 3.8 hold, let $\mathcal{Z}_{0}$ be any compact set of $\mathbb{R}^{m}$ and ${\bm{Z}}_{0}={\bm{z}}_{0}{\bm{1}}^{\top},{\bm{z}}_{0}\in\mathcal{Z}_{0}$. $\{{\bm{Z}}_{k}\}$ is generated by (DSM) and the evaluation noise is introduced by random reshuffling (3.8). Then for any $\hat{\varepsilon}\in(0,1)$ and $M>0$, there exists a constant $\alpha>0$ such that for any $k\geq 0$, any $\eta_{kN}\in(0,\alpha]$, we have
@@
Therefore, for any $k>0$ and any $10\leq l\leq N-$ 1, the following inclusion holds:
@@
[…]
@@
Lemma 3.9 illustrates that $\{\frac{1}{d}{\bm{Z}}_{kN\land\tau^{\prime}_{M}}{\bm{1}}_{d}\}$ can be viewed as an inexact discretization of differential inclusion (3.1). Then based on Lemma 2.15, we establish the uniform boundedness and asymptotic convergence of $\{{\bm{Z}}_{k}\}$ when adopting (3.8) under Assumption 3.8. As the evaluation noise in (3.9) vanishes over each epoch and is thus asymptotically controlled, Theorems 3.10 and 3.11 can be viewed as special cases of Theorems 3.5 and 3.7.
@@
Let $\mathcal{Z}_{0}$ be any compact set of $\mathbb{R}^{m}$. Suppose Assumption 3.1 and 3.8 hold, and ${\bm{Z}}_{0}={\bm{z}}_{0}{\bm{1}}^{\top}_{d},$ where ${\bm{z}}_{0}\in\mathcal{Z}_{0}$. Then for any $r>\max\{0,4\sup_{{\bm{z}}\in\mathcal{Z}_{0}\cup\mathcal{A}}\psi({\bm{z}})\}$, there exist $\alpha>0$, such that for any $\alpha$-upper-bounded $\{\eta_{k}\}$, the sequence $\{{\bm{Z}}_{k}\}$ generated by (DSM) with (3.8) is restricted in $\mathcal{L}_{2r}^{d}$.
@@
[…]
@@
[…]
@@ 3.3 Convergence under with-replacement sampling
Suppose Assumption 3.1 and 3.12 hold. Then for any $\varepsilon,T,\iota>0$, there exists $\beta_{0}>0$, such that for any $\beta_{0}$-upper-bounded sequence $\{\eta_{k}\}$, it holds that
($\mathbb{P}\left(\left\{\exists s\geq 0,\text{ such that }\sup_{s\leq i\leq\Lambda_{\eta}(\lambda_{\eta}(s)+T)}\left\|\sum_{k=s}^{i}\eta_{k}\frac{1}{d}\Xi_{k+1}{\bm{1}}_{d}\mathbb{1}_{\tau_{M}>k}\right\|\geq\iota\right\}\right)\leq\varepsilon.$
Lemma 3.13 is a direct corollary of [46, Proposition 4.2] and [43, Proposition A.5], hence we omit its proof for simplicity. Based on this lemma, we present Theorem 3.14 and 3.15 to show the uniform boundedness and convergence properties of framework (DSM) together with evaluation noise (3.18).
@@
Suppose Assumption 3.1 and 3.12 hold, and let $\mathcal{Z}_{0}$ be any compact set of $\mathbb{R}^{m}$, ${\bm{Z}}_{0}={\bm{z}}_{0}{\bm{1}}_{d}^{\top},{\bm{z}}_{0}\in\mathcal{Z}_{0}$. The sequence $\{{\bm{Z}}_{k}\}$ is generated by (DSM), and the evaluation noise is introduced by with-replacement sampling (3.18). Then for any $r>\max\{0,4\sup_{x\in\mathcal{Z}_{0}\cup\mathcal{A}}\psi(x)\}$ and any $\varepsilon\in(0,1)$, there exists $\beta>0$, such that for any $\beta$-upper-bounded sequence $\{\eta_{k}\}$ and $\{\delta_{k}\}$, $\{{\bm{Z}}_{k}\}$ is restricted in $\mathcal{L}_{2r}^{d}$ with probability at least $1-\varepsilon$.
@@
Suppose Assumption 3.1 and 3.12 hold, $\{\delta_{k}\}$ in (1.5) satisfies $\lim_{k\to\infty}\delta_{k}=0$. Let $\mathcal{Z}_{0}$ be any compact set of $\mathbb{R}^{m}$, and ${\bm{Z}}_{0}={\bm{z}}_{0}{\bm{1}}_{d}^{\top},{\bm{z}}_{0}\in\mathcal{Z}_{0}$. Then for any $\varepsilon\in(0,1)$, there exists $\beta>0$, such that for any $\{\eta_{k}\},\{\delta_{k}\}$ upper-bounded by $\beta$, and any $\{{\bm{Z}}_{k}\}$ generated by (DSM) with evaluation noise (3.18), we have
@@
Denote the event $\Omega_{1}:=\{\{{\bm{Z}}_{k}\}\subseteq\mathcal{L}_{2r}^{d}\}$. By Theorem 3.5 and 3.14, we have
@@
Notice that $\{\frac{1}{d}\Xi_{k+1}{\bm{1}}_{d}\}$ is a uniformly bounded martingale difference sequence over $\Omega_{1}$ and $\eta_{k}\sim o(1/\log(k))$. Invoking [45, Proposition 1.4, Remark 1.5], we obtain that

### v4

@@ 2.2 Mixing matrix
Mixing matrix ${\bm{W}}=[{\bm{W}}(i,j)]\in\mathbb{R}^{d\times d}$ determines the structure of the communication network and plays an important role in averaging the local information of neighbor agents. Generally, we assume the mixing matrix $\bm{W}$} satisfies the following properties, which are standard in the literature [xiao2004fast, sundhar2010distributed, lian2017can].
@@
With a given graph $\mathtt{G}$, there are various approaches to choose its corresponding mixing matrix, such as Laplacian-based constant edge weight matrix [xiao2004fast] and Metropolis weight matrix [xiao2006distributed]. We refer interested readers to [nedic2018network, shi2015extra, wang2022decentralized, wang2025decentralized] for more details about how to choose the mixing matrix.
Based on [pillai2005perron, Perron-Frobenius Theorem], we derive a direct corollary, which illustrates the fundamental spectral property of mixing matrix $\bm{W}$}.
@@
For any mixing matrix ${\bm{W}}\in\mathbb{R}^{d\times d}$ that corresponds to a connected graph $\mathtt{G}$, all the eigenvalues of $\bm{W}$} lie in (1,1](-1,1], and $\bm{W}$} has a single eigenvalue at $1$ that admits $\bm{1}_{d}$} as its eigenvector.
@@ 2.3 Nonsmooth analysis and conservative field
[clarke1990optimization] For any given locally Lipschitz continuous function $f:\mathbb{R}^{n}\to\mathbb{R}$, and any ${\bm{x}}\in\mathbb{R}^{n}$, the generalized directional derivative of $f$ at $\bm{x}$} along the direction ${\bm{d}}\in\mathbb{R}^{n}$, is defined as
@@
[clarke1990optimization] We say that $f$ is Clarke regular at ${\bm{x}}\in\mathbb{R}^{n}$, if for every direction ${\bm{d}}\in\mathbb{R}^{n}$, the one-sided directional derivative
@@
It is remarkable to note that any conservative field is locally bounded [bolte2021conservative, Remark 3]. We now introduce the definition of the path-differentiable function corresponding to the conservative field.
@@
As illustrated in [bolte2021conservative, Corollary 1], for any given path-differentiable function $f$, $\partial f$ is a conservative field for $f$. Therefore, the concept of the conservative field can be regarded as an extension of the Clarke subdifferential. More importantly, when conservative field $D_{f}({\bm{x}})$ is convex-valued, ${\bm{0}}\in\partial f({\bm{x}})$ implies that $0\in D_{f}({\bm{x}})$.
In fact, the class of path-differentiable functions is general enough to cover the most objective functions in real-world problems. The well-known Clarke regular functions mentioned in [clarke1990optimization] and semi-algebraic functions mentioned in [ojasiewicz1965EnsemblesS] are all path-differentiable functions. In [davis2020stochastic, bolte2021conservative], the authors identify definable functions (whose graphs are definable in an o-minimal structure) as an important class of path-differentiable functions. Actually, most activation functions and loss functions in deep neural networks are definable functions, including sigmoid, softplus, ReLU, leaky ReLU, hinge loss, etc.
Furthermore, it is important to highlight that definability is preserved under finite summation and composition [davis2020stochastic]. Consequently, for any neural network constructed by definable blocks, its loss function is definable, thus it is a path-differentiable function. Additionally, it is worth noting that the Clarke subdifferential of definable functions is itself definable [bolte2021conservative]. Hence, in the case of a neural network constructed by definable blocks, the conservative field corresponding to the AD algorithms is also definable. The following proposition demonstrates that the definability of both $f$ and its conservative field $D_{f}$ leads to the nonsmooth Morse–Sard property [bolte2007clarke].
Proposition 2.7 (Theorem 5 in [bolte2021conservative]).
@@ 2.4 Differential inclusion and stochastic subgradient methods
Lemma 2.13 (Extension of Proposition 1.3 in [benaim2005stochastic]).
@@
Before exhibiting the results on uniform boundedness and convergence of (2.3), we provide an assumption about differential inclusion (2.2), which is commonly used in the analysis of ODE-based approaches [benaim2006dynamics, borkar2009stochastic, duchi2018stochastic].
@@
Lemma 2.15 (Theorem 3.6 in [xiao2023convergence]).
@@
Lemma 2.16 (Summary of Theorem 3.6, Proposition 3.27 in [benaim2005stochastic]).
@@
Except for Lemma 2.16, similar results under slightly different conditions can be found in [borkar2009stochastic, davis2020stochastic, duchi2018stochastic], while some recent works [bianchi2021closed, bolte2022long] focus on analyzing the convergence of (2.3) under relaxed conditions. Interested readers could refer to those works for details.
@@ 3 A General Framework for Decentralized Stochastic Subgradient-based Methods
In this section, we aim to establish the convergence properties of proposed framework (DSM) under two noise settings. To this end, it is crucial to ensure the uniform boundedness of iterates $\{{\bm{Z}}_{k}\}$. A large amount of literature [castera2021inertial, ruszczynski2020convergence, davis2020stochastic, le2024nonsmooth] regards this condition as a prior assumption when analyzing the convergence of SGD-type algorithms in a single-agent setting. Nevertheless, more recent works [bianchi2019constant, bianchi2022convergence, josz2024global, josz2023lyapunov, bolte2024inexact, xiao2023convergence] focus on rigorously establishing the uniform boundedness of the iterative sequence.
@@ 3.1 Basic assumptions and main results
Assumption 3.1(1) allows for a flexible choice of the step-size, while Assumption 3.1(2) is reasonable and mild, which is commonly employed in various existing works [bolte2021conservative, davis2020stochastic, xiao2023convergence, xiao2023adam, castera2021inertial].
Assumption 3.1(3) clarifies the Lyapunov properties of a differential inclusion with respect to $\Phi$, which is standard in the literature [benaim2005stochastic, borkar2008stochastic, bian2009subgradient, davis2020stochastic, bolte2021conservative, josz2023lyapunov] and can be satisfied by most appropriately selected set-valued mappings. Since each $\Phi_{i}$ is graph-closed and locally bounded, the convex-valued set-valued mapping $\Phi$ is also graph-closed and locally bounded. Then, the coercivity of $\psi$ in Assumption 3.1(3) indicates the compactness of any level set $\mathcal{L}_{r}:=\{x\in\mathbb{R}^{n}:\psi(x)\leq r\}$.
Assumption 3.1(4) is similar to the Weak Sard’s condition in [davis2020stochastic], which stipulates that the set of values of $\psi$ at points outside the stable set is dense in $\mathbb{R}$. As demonstrated in [bolte2021conservative], Assumption 3.1(4) holds when $f$ is a definable function with the selection $\Phi:=\partial f$ and $\psi:=f$.
@@
$\leq{}$ $\left\|{\bm{Z}}_{k\land(\tau_{M}-1)}{\bm{W}}({\bm{I}}_{d}-{\bm{P}})\right\|+\left\|\eta_{k\land(\tau_{M}-1)}({\bm{H}}_{k\land(\tau_{M}-1)}+\Xi_{(k+1)\land\tau_{M}})({\bm{I}}_{d}-{\bm{P}})\right\|$
$\leq{}$ $\left\|{\bm{Z}}_{k\land(\tau_{M}-1)}({\bm{I}}_{d}-{\bm{P}})\right\|\left\|{\bm{W}}({\bm{I}}_{d}-{\bm{P}})\right\|+\left\|\eta_{k\land(\tau_{M}-1)}({\bm{H}}_{k\land(\tau_{M}-1)}+\Xi_{(k+1)\land\tau_{M}})\right\|$
$\leq{}$ $\lambda_{2}\left\|{\bm{Z}}_{\perp,k\land(\tau_{M}-1)}\right\|+\left\|\eta_{k\land(\tau_{M}-1)}({\bm{H}}_{k\land(\tau_{M}-1)}+\Xi_{(k+1)\land\tau_{M}})\right\|$
$\leq{}$ $\sum_{i=0}^{k\land(\tau_{M}-1)}\lambda_{2}^{k\land(\tau_{M}-1)-i}\|\eta_{i}({\bm{H}}_{i}+\Xi_{i+1})\|+\lambda_{2}^{(k+1)\land\tau_{M}}\|{\bm{Z}}_{\perp,0}\|$
$\leq{}$ $\frac{\sup_{0\leq i\leq k\land(\tau_{M}-1)}\eta_{i}\ell_{M}}{1-\lambda_{2}},$
@@
Remark 3.5.
It is worth mentioning that the value of $\lambda_{2}$ (i.e., the second-largest eigenvalue of the mixing matrix $\bm{W}$}) is a critical factor in quantifying the effect of network topology on the convergence of decentralized methods. As demonstrated in [boyd2004fastest], $\lambda_{2}$ is widely regarded as a measurement for network connectivity. Proposition 3.4 establishes (3.6), demonstrating that $\lambda_{2}$ governs the rate at which $\bm{Z}_{\perp,k}$ converges to zero. Consequently, the value of $\lambda_{2}$ determines how rapidly the iterates $\frac{1}{d}{\bm{Z}}_{(k+1)\land\tau_{M}}{\bm{1}}_{d}$} approach the trajectory of the differential inclusion (3.1). For further analysis regarding the influence of network topology in smooth settings, we refer readers to some related works [vogels2023beyond, song2022communication].
@@
The following lemma from [sundhar2010distributed, Lemma 3.1] concerns the limit property of a convolution-like scalar sequence.
@@
Let $\{\gamma_{k}\}$ be a sequence of real numbers. If $0<a<1$ and $\lim_{k\rightarrow\infty}\gamma_{k}=\gamma$, then it holds that $\lim_{k\rightarrow\infty}\sum_{\ell=0}^{k}a^{k-\ell}\gamma_{\ell}=\frac{\gamma}{1-a}$.
@@ 3.2 Convergence with random reshuffling
Suppose Assumption 3.1 and 3.9 hold, let $\mathcal{Z}_{0}$ be any compact set of $\mathbb{R}^{m}$ and ${\bm{Z}}_{0}={\bm{z}}_{0}{\bm{1}}^{\top},{\bm{z}}_{0}\in\mathcal{Z}_{0}$. $\{{\bm{Z}}_{k}\}$ is generated by (DSM) and the evaluation noise is introduced by random reshuffling (3.8). Then for any $\hat{\varepsilon}\in(0,1)$ and $M>0$, there exists a constant $\alpha>0$ such that for any $k\geq 0$, any $\eta_{kN}\in(0,\alpha]$, we have
@@
Therefore, for any $k>0$ and any $0\leq l\leq N-1$, the following inclusion holds:
@@
$\in{}$ $\frac{1}{d}{\bm{Z}}_{kN\land\tau^{\prime}_{M}}{\bm{1}}_{d}-\sum_{l=kN}^{(k+1)N-1}\eta_{kN}\frac{1}{d}\sum_{i=1}^{d}\left[\mathcal{U}_{i,i_{l}}(\frac{1}{d}{\bm{Z}}_{kN\land\tau^{\prime}_{M}}{\bm{1}}_{d}+\mathbb{B}({\bm{0}},\hat{\varepsilon}))+\mathbb{B}({\bm{0}},\frac{\hat{\varepsilon}}{3})\right]\mathbb{1}_{\tau^{\prime}_{M}>kN}$
$\subseteq{}$ $\frac{1}{d}{\bm{Z}}_{kN\land\tau^{\prime}_{M}}{\bm{1}}_{d}-N\eta_{kN}\left(\frac{1}{d}\sum_{i=1}^{d}\frac{1}{N}\sum_{l=0}^{N-1}\mathcal{U}_{i,l}(\frac{1}{d}{\bm{Z}}_{kN\land\tau^{\prime}_{M}}{\bm{1}}_{d}+\mathbb{B}({\bm{0}},\hat{\varepsilon}))+\mathbb{B}({\bm{0}},\frac{\hat{\varepsilon}}{3})\right)\mathbb{1}_{\tau^{\prime}_{M}>kN}$
$\subseteq{}$ $\frac{1}{d}{\bm{Z}}_{kN\land\tau^{\prime}_{M}}{\bm{1}}_{d}-N\eta_{kN}\Phi^{\hat{\varepsilon}}\left(\frac{1}{d}{\bm{Z}}_{kN\land\tau^{\prime}_{M}}{\bm{1}}_{d}\right)\mathbb{1}_{\tau^{\prime}_{M}>kN},$
@@
Lemma 3.10 illustrates that $\{\frac{1}{d}{\bm{Z}}_{kN\land\tau^{\prime}_{M}}{\bm{1}}_{d}\}$ can be viewed as an inexact discretization of differential inclusion (3.1). Then based on Lemma 2.15, we establish the uniform boundedness and asymptotic convergence of $\{{\bm{Z}}_{k}\}$ when adopting (3.8) under Assumption 3.9. As the evaluation noise in (3.9) vanishes over each epoch and is thus asymptotically controlled, Theorems 3.11 and 3.12 can be viewed as special cases of Theorems 3.6 and 3.8.
@@
Let $\mathcal{Z}_{0}$ be any compact set of $\mathbb{R}^{m}$. Suppose Assumption 3.1 and 3.9 hold, and ${\bm{Z}}_{0}={\bm{z}}_{0}{\bm{1}}^{\top}_{d},$ where ${\bm{z}}_{0}\in\mathcal{Z}_{0}$. Then for any $r>\max\{0,4\sup_{{\bm{z}}\in\mathcal{Z}_{0}\cup\mathcal{A}}\psi({\bm{z}})\}$, there exist $\alpha>0$, such that for any $\alpha$-upper-bounded $\{\eta_{k}\}$, the sequence $\{{\bm{Z}}_{k}\}$ generated by (DSM) with (3.8) is restricted in $\mathcal{L}_{2r}^{d}$.
@@
$\leq{}$ $\left\|{\bm{z}}_{i,(kN+j)\land\tau_{M}}-\frac{1}{d}{\bm{Z}}_{(kN+j)\land\tau_{M}}{\bm{1}}_{d}\right\|+\left\|\frac{1}{d}{\bm{Z}}_{(kN+j)\land\tau_{M}}{\bm{1}}_{d}-\frac{1}{d}{\bm{Z}}_{kN}{\bm{1}}_{d}\right\|$
$<{}$ $\frac{2\alpha_{1}}{3}<\frac{r}{L_{2}}.$
@@
Suppose Assumption 3.1 and 3.9 hold. Let $\mathcal{Z}_{0}$ be any compact set of $\mathbb{R}^{m}$, $\{{\bm{Z}}_{k}\}$ be generated by (DSM) with (3.8), and ${\bm{Z}}_{0}={\bm{z}}_{0}{\bm{1}}^{\top}_{d},{\bm{z}}_{0}\in\mathcal{Z}_{0}$. Then there exists $\alpha_{\text{rr}}>0$, such that for any $\alpha_{\text{rr}}$-upper-bounded diminishing $\{\eta_{k}\}$ in (DSM), it follows that
@@ 3.3 Convergence under with-replacement sampling
Suppose Assumption 3.1 and 3.13 hold. Then for any $\varepsilon,T,\iota>0$, there exists $\beta_{0}>0$, such that for any $\beta_{0}$-upper-bounded sequence $\{\eta_{k}\}$, it holds that
$\mathbb{P}\left(\left\{\exists s\geq 0,\text{ such that }\sup_{s\leq i\leq\Lambda_{\eta}(\lambda_{\eta}(s)+T)}\left\|\sum_{k=s}^{i}\eta_{k}\frac{1}{d}\Xi_{k+1}{\bm{1}}_{d}\mathbb{1}_{\tau_{M}>k}\right\|\geq\iota\right\}\right)\leq\varepsilon.$
Lemma 3.14 is a direct corollary of [benaim2006dynamics, Proposition 4.2] and [xiao2023convergence, Proposition A.5], hence we omit its proof for simplicity. Based on this lemma, we present Theorem 3.15 and 3.16 to show the uniform boundedness and convergence properties of framework (DSM) together with evaluation noise (3.18).
@@
Suppose Assumption 3.1 and 3.13 hold, and let $\mathcal{Z}_{0}$ be any compact set of $\mathbb{R}^{m}$, ${\bm{Z}}_{0}={\bm{z}}_{0}{\bm{1}}_{d}^{\top},{\bm{z}}_{0}\in\mathcal{Z}_{0}$. The sequence $\{{\bm{Z}}_{k}\}$ is generated by (DSM), and the evaluation noise is introduced by with-replacement sampling (3.18). Then for any $r>\max\{0,4\sup_{x\in\mathcal{Z}_{0}\cup\mathcal{A}}\psi(x)\}$ and any $\varepsilon\in(0,1)$, there exists $\beta>0$, such that for any $\beta$-upper-bounded sequence $\{\eta_{k}\}$ and $\{\delta_{k}\}$, $\{{\bm{Z}}_{k}\}$ is restricted in $\mathcal{L}_{2r}^{d}$ with probability at least $1-\varepsilon$.
@@
Suppose Assumption 3.1 and 3.13 hold, $\{\delta_{k}\}$ in (1.5) satisfies $\lim_{k\to\infty}\delta_{k}=0$. Let $\mathcal{Z}_{0}$ be any compact set of $\mathbb{R}^{m}$, and ${\bm{Z}}_{0}={\bm{z}}_{0}{\bm{1}}_{d}^{\top},{\bm{z}}_{0}\in\mathcal{Z}_{0}$. Then for any $\varepsilon\in(0,1)$, there exists $\beta>0$, such that for any $\{\eta_{k}\},\{\delta_{k}\}$ upper-bounded by $\beta$, and any $\{{\bm{Z}}_{k}\}$ generated by (DSM) with evaluation noise (3.18), we have
@@
Denote the event $\Omega_{1}:=\{\{{\bm{Z}}_{k}\}\subseteq\mathcal{L}_{2r}^{d}\}$. By Theorem 3.6 and 3.15, we have
@@
Notice that $\{\frac{1}{d}\Xi_{k+1}{\bm{1}}_{d}\}$ is a uniformly bounded martingale difference sequence over $\Omega_{1}$ and $\eta_{k}\sim o(1/\log(k))$. Invoking [benaim2005stochastic, Proposition 1.4, Remark 1.5], we obtain that